OpenStudy (anonymous):

Solve: -2t + 7 – 5t ≥ 3 + 4t + 10. A. t ≤ -20/7 B. t ≤ -6/11 C. t ≥ -20/7 D. t ≥ -6/11

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (johnweldon1993):

Hello again :) \[\large -2t + 7 - 5t \ge 3 + 4t + 10\] We need to combine like terms here...look how we have \(\large -2t - 5t\) on the left hand side...and notice how we have 3 + 10 on the right hand side...lets simplify that...

OpenStudy (anonymous):

-7t and 13

OpenStudy (johnweldon1993):

Right...so we have \[\large -7t + 7 \ge 4t + 13\] Now we want o continue to combine like terms..but by getting the like terms on 1 side... So we want all the terms with a 't' on 1 side...and all the ones without..on the other.. So lets subtract 7 from both sides of this equation...and then subtract 4t from both sides...what do we have then?

OpenStudy (anonymous):

7>13

OpenStudy (anonymous):

wait

OpenStudy (johnweldon1993):

I'm waiting ... :)

OpenStudy (anonymous):

idk im lost i know i did something wrong

OpenStudy (anonymous):

ok explain you take 7 form both sides frist right

OpenStudy (johnweldon1993):

Here I already typed out the explanation :) Not quite...so we would have \[\large -7t + 7 \ge 4t + 13\] We want to subtract 7 from both sides of the equation \[\large -7t + 7 \color \red{- 7} \ge 4t + 13 \color \red{-7}\] which will give us \[\large -7t \ge 4t + 6\] Now we want to subtract 4t from both sides of the equation \[\large -7t \color \red{-4t}\ge 4t \color \red{-4t} + 6\] Which will then give us \[\large -11t \ge 6\]

OpenStudy (anonymous):

but then thats the answer which one is it ?

OpenStudy (johnweldon1993):

That isn't the answer...remember we want to solve for 't'...not -11t So we need to divide both sides of that equation by -11 \[\large t \le -\frac{6}{11}\] ****************notice what happened with the sign there....******************* We HAD \(\ge\) but we ended with \(\le\) this is because we divided by a negative number...when that happens we switch the direction of the sign

OpenStudy (anonymous):

Your lots of help i really like how your explaining everything

OpenStudy (johnweldon1993):

Glad to help :)